Grade 1110 grade maths

Q) A string of length L is fixed at one end and carries a mass M at the other end. The string makes 2/pi revolution per second around the vertical axis through the fixed end. The tension in the string is?

Yash Badgujar
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
Dear Yash
length of string = L
mass = M
angular speed (w) = 2/ pi    revolutions/sec  = 2*2 pi / pi = 4 rad/sec
tension (T) = m *v2 / r
 = m* r* w2
 = M * L * 42
 = 16 ML Newton
Regards
Arun (askIITians forum expert)